any clue anyone knows how to answer any of these or attempt to?


Q3. The Trypsin-EDTA solution used in the experiment would contain 0.4% trypsin (w/v) and 

0.03% EDTA (w/v). It is provided to you as a 10x stock solution.   

(a)  EDTA has a molecular weight of 292.24. What is the molarity of the EDTA in the 10X stock solution? 

(b)  What is the total final amount (in mg) of Trypsin which is added to the cells?

 

Q4. Following a 2 minute incubation, FBS solution is then added to the cells to deactivate the trypsin enzyme. The solution comprises 10% FBS, made up in 1x PBS. You decide to make up 50 ml of this solution using your 10 x PBS stock, and a 100% FBS stock. 

(a) How much 10 x PBS will you need to use to make the 50 ml solution?

(b) How much 100% stock FBS would you need?




1
Expert's answer
2021-03-23T10:14:04-0400



Q3

(a) The molarity can be calculated using the equation:

M = n / V = m / MrV

where M - molarity, n - number of moles, V - volume, m - mass, Mr - molecular weight.

As 10x solution contains 0.03% EDTA, 100 mL of the solution contain 0.03 g of EDTA. From here:

M(EDTA) = 0.03 g / (292.24 g/mol × 0.1 L) = 0. mol/L = 1.027 mM

(b) If 1 mL of solution is added to the cells, the final amount of trypsin is:

m = (0.4 g / 100) / 10 = 0.0004 g = 0.4 mg


Q4

(a) For preparing 50 mL of 10% FBS in 1 x PBS:

V(PBS) = 50 mL / 10 = 5 mL

As a result, 5 mL of 10 x PBS are needed.

(b) For preparing 50 mL of 10% FBS in 1 x PBS:

V(FBS) = (50 mL / 100%) × 10% = 5 mL

As a result, 5 mL of 100% FBS are needed.


Answer: Q3: (a) 1.027 mM, (b) 0.4 mg; Q4: 5 mL; 5 mL.